An air column which has a fixed length 0.78 m is closed at one end and open at the other end. It produces a 2nd resonant frequency of 335 Hz. Calculate the speed of sound in air.


1
Expert's answer
2021-02-03T02:47:49-0500

A closed at one end tube will have approximate resonances of:


"f = \dfrac{nv}{4L}"

where "n" is an odd number (order of resonant frequency), "v" is the speed of sound, and "L = 0.78m" is the length of the tube. Since "f = 335Hz", and it is the 2nd resonant frequency the corresponding order will be "n=3". Thus, obtain the expression for the speed of sound:


"v = \dfrac{4fL}{n}\\
v = \dfrac{4\cdot 335\cdot 0.78}{3} = 348.4m/s"

Answer. 348.4 m/s.